OpenStudy (anonymous):

help me simplify these complex expressions? :) Algebra 2

OpenStudy (anonymous):

ANSWER CHOICES: 7i over 10, -7i over 10, -7 over 10, 70i over 100i^2.

OpenStudy (anonymous):

\[6\sqrt{3} , -6\sqrt{3} , 6i \sqrt{3} , -6i \sqrt{3}\]

OpenStudy (anonymous):

technically, i is a radical since \(\large i=\sqrt{-1} \), so to simplifiy \(\large \frac{7}{10i} \), you would multiply both numerator and denominator by i....

OpenStudy (anonymous):

thankyou, but i'm not quite sure what that would be. 7*sqrt -1 divided by 10i* sqrt -1 ?? @dpaInc

OpenStudy (anonymous):

\[\huge \frac{7}{10i} = \frac{7}{10i} \cdot \frac{i}{i}= \] can you simplify it from here?

OpenStudy (anonymous):

ie, simplifying complex numbers is like simplifying radicals....

OpenStudy (anonymous):

well, the only thing i know to do here is cross mulitply but you can't do that @dpaInc

OpenStudy (anonymous):

cross multiply? you're multiplying fractions here.... just go straight across....

OpenStudy (anonymous):

what the numerator gonna be? denominator?

OpenStudy (anonymous):

14i over 20 i?

OpenStudy (anonymous):

how did you get 14i for the numerator? 7 times i = 7i...

OpenStudy (anonymous):

and the denominator: 10i times i = \(\large 10i^2 \)

OpenStudy (anonymous):

oh goodness. :0